Write me Heoderson-Hasselbalch equation for a prop
Answer #1

Henderson-Hasselbalch:

pH=pKa+log([A-]/[HA])

log([A-]/[HA]) = pH-pKa

log([A-]/[HA]) = pH - 4.874

[A-]/[HA]=10pH - 4.874

[A-]/[HA]=104.550 - 4.874=0.47

[A-]/[HA]=105.407 - 4.874=3.41

[A-]/[HA]=104.874 - 4.874=100=1